Blossom, Inc., is considering investing in a new production line for eye drops. Other than investing in the equipment, the company needs to increase its cash and cash equivalents by $10,000, increase the level of inventory by $25,000, increase accounts receivable by $25,000, and increase accounts payable by $5,000 at the beginning of the project. Blossom will recover these changes in working capital at the end of the project 14 years later. Assume the appropriate discount rate is 12 percent. What are the present values of the relevant investment cash flows?
Answer #1

INVESTMENT IN WORKING CAPITAL = 10000 +25000 + 25000 - 5000 = 55000

THE ABOVE AMOUNT RECOVERED AT THE END OF 14 YEARS

SO PRESENT VALUE = WORKING CAPITAL/(1+r)^n = 55000/(1+0.12)^14 = 11254.09 (rounded to 2 decimals)

Answer : 11254.09 (Thumbs up please)